# Supplement

### v3

@@ 2.5 Alignment, Estimation, and Convergence
To monitor convergence of FAST we use the Gelman-Rubin RHat statistic for all model parameters (Gelman and Rubin,, 1992), including the FPCs. This is done after post-processing the PCs and scores.
@@ 3 Bayesian implementation in STAN
In this section, we provide the simple STAN Carpenter et al., (2017) implementation of FAST. We begin with the data section, which declares constants and imports the data, spline basis, and penalty matrix. All notation is as used in Section 2.
@@
int M; // Number of observations int Q; // Number of spline bases int K; // Number of Eigenfunctions
matrix[N, M] Y; // Functional data matrix[M, Q] B; // Orthogonal spline basis matrix[Q, Q] P; // Penalty matrix for splines }
In the parameters block, we model the matrix $\mathbf{X}$ (X) rather than $\mathbf{\Psi}$, which is obtained from $\mathbf{X}$ via polar decomposition. This block also contains the eigenvalues $\lambda_{k}$ (lambda) stored using a positive_ordered type, ordering the eigenvalues and corresponding eigenfunctions. The matrix Scores is the $N\times K$ score matrix $\mathbf{\Xi}$ introduced in Result 1.
@@
// Mean structure vector[Q] w_mu; // Population mean parameters real<lower=0> h_mu; // Population mean smoothing parameter
// Covariance structure positive_ordered[K] lambda; // Eigenvalues vector<lower=0>[K] H; // EF smoothing parameters matrix[Q, K] X; // Unconstrained EF weights (X) matrix[N, K] Scores; // EF scores (xi) }
@@
// Polar decomposition matrix[Q,K] Psi; { matrix[K,K] evec_XtX = eigenvectors_sym(X’*X); vector[K] eval_XtX = eigenvalues_sym(X’*X); Psi = X*evec_XtX*diag_matrix(1/sqrt(eval_XtX))*evec_XtX’; } }
The model proceeds with the specification of priors within the model section. The key component is the prior on $\mathbf{\Psi}$ (Psi), which just requires specifying independent $N(0,1)$ priors for each entry of the matrix $\mathbf{X}$ (X). Combined with the polar decomposition in the transformed parameters block, this induces a uniform prior on $\mathbf{\Psi}$ over the Stiefel manifold, $\mathcal{V}_{K,Q}$. FAST thus models the FPCs through the small-dimensional $\mathbf{X}\in\mathbb{R}^{Q\times K}$, constrained by the size of the spline basis and number of FPCs. This dimensionality reduction likely contributes to the much improved computational efficiency. All other priors are as described in Section 2.3.
@@
H $\sim$ gamma(0.001, 0.001); h_mu $\sim$ gamma(0.001, 0.001); lambda $\sim$ inv_gamma(0.001, 0.001); sigma2 $\sim$ inv_gamma(0.001, 0.001);
// Smoothing additions to the target density target += Q/2.0 * log(h_mu) - h_mu / 2.0 * w_mu’ * P * w_mu; for(i in 1:K){ target += Q/2.0 * log(H[i]) - H[i] / 2.0 * Psi[,i]’ * P * Psi[,i]; }
// Indirect uniform prior on Stiefel manifold to_vector(X) $\sim$ normal(0,1);
[…]
@@
model{
[…]
@@ 4 Simulations
We consider the following two simulation scenarios for model (1), where $f_{k}(t)$ is are the Legendre polynomials on $[0,1]$:
@@
For each scenario we generate $200$ datasets with $N=50$ functions observed at $M=50$ points distributed along $[0,1]$ according to quadrature. FAST is compared to the Generalized Function-on-Scalar Regression of Goldsmith et al., (2015) (labeled “GFSR”), polar expansion in the observed data space from Jauch et al., (2021) (labeled “POLAR”), and Variational Message Passing from Nolan et al., (2023) (labeled “VMP”). We do not compare with Ye, (2024) due to there being no available software. For the MCMC-based approaches, we performed 1500 sampling iterations and discarded the first 1000 as burn-in. This was chosen based upon the convergence properties of FAST and kept uniform between methods.
To obtain the orthonormal FPCs from the unconstrained samples of GFSR, we leverage the orthogonalization procedure suggested by Goldsmith et al., (2015). This involves calculating the SVD of the unconstrained FPC matrix sample and taking the right singular vectors. The scores are correspondingly transformed by the left singular vectors and diagonal matrix of singular values.
As FAST, POLAR, and GFSR all have the same identifiability issues described in Section 2.5, we applied the post-processing steps described in that section to these three methods. FPC/score signs were then aligned with those of the true underlying FPCs for the purposes of evaluation. Similar to FAST, both POLAR and GFSR require procedures to produce orthonormal (in the vector sense) FPC estimates. For POLAR, Jauch et al., (2021) recommends taking the right singular vectors of the mean smoothed data. As for GFSR, we take the point-wise mean of the non-orthogonal latent principal component matrix, then apply the SVD-based rotation Goldsmith et al., (2015) describes. VMP does not provide explicit samples, just FPC estimates.
POLAR, GFSR, and VMP all produce FPC estimates/samples which are orthonormal in the vector sense. So, we rescale the FPC estimates for these methods to have norm 1 in $L^{2}([0,1])$, estimated using Gaussian quadrature weights.
@@
We also estimate pointwise posterior coverage probabilities of the true FPCs $\phi_{k}(t)$, $k=1,\ldots,K$. For each simulated data set $b\in\{1,\ldots,B\}$, we obtain the equal-tail $95$% credible intervals at the observed time points $t_{m}$, $m=1,\ldots,M$. The proportion of times these credible intervals cover the true function is calculated over the $t_{m}$ for each combination of $b$ and $k$. These proportions are then collected in a vector of length $B=200$ for each function $\phi_{k}(t)$. VMP does not produce credible intervals for the FPCs, so it is excluded from this comparison.
Figure 2: Kernel smoother of $95$% credible interval coverage probabilities of the true FPCs for FAST, POLAR, and GFSR. Row indicates simulation scenario, S1 then S2. Each column corresponds to a particular eigenfunction. Distribution means: horizontal solid lines; nominal $95$% level: horizontal dotted lines.
Figure 2 displays the kernel smooth of the estimated point-wise coverage proportions for FAST, POLAR, and GFSR using the same structure as Figure 1. FAST performs well in all scenarios, consistently out-performing the other approaches.
Figures 1 and  2 indicate that FAST has an advantage in all scenarios, but the advantage is particularly striking for $\phi_{3}(t)$ in S1. In this case, the mean can be written as a linear combination of $\phi_{1}(t)$ and $\phi_{3}(t)$: $\mu(t)=140\phi_{1}(t)+4\sqrt{5}\phi_{3}(t)$. This choice makes the problem more challenging in finite samples, when information may leak between the mean and covariance; see, example, Zhou et al., (2025) for a more detailed discussion.
@@
We also compared estimation accuracy and posterior coverage of credible intervals for the scores $\xi_{k}$. Accuracy is calculated as mean square error aggregated by FPC, $k$, within each simulation, $b$. Coverage was calculated as the proportion of times the equal-tail $95$% credible intervals cover the corresponding true scores; see results in Supplement Section S3.3. FAST outperforms the other methods and has near nominal coverage.
@@
$M$ FAST GFSR POLAR VMP $N$ FAST GFSR POLAR VMP
50 $2.9$ $24.3$ $9.2$ $0.2$ 50 $2.9$ $24.3$ $9.2$ $0.2$
100 $5.9$ $39.5$ $28.5$ $1.0$ 100 $3.5$ $43.2$ $15.4$ $0.6$
250 $13.0$ $82.8$ $1.2$ 250 $5.8$ $47.2$ $29.0$ $0.6$
500 $37.3$ $151.0$ $5.2$ 500 $14.4$ $27.8$ $52.1$ $0.7$
@@
FAST and GFSR scale approximately linearly with $M$, though GFSR has a substantially higher slope. For $N$, FAST and POLAR scale linearly, while GFSR seems to be more performant with higher $N$, possibly due to the increased information better specifying the posterior geometry. We could not fit POLAR for $M\geq 250$ because it uses too much memory. VMP is the fastest, not unexpected given the approximation inherent to the technique. All MCMC-based methods were run for 1500 iterations with the first 1000 discarded as burn-in; however, the convergence characteristics were not uniform. FAST and POLAR exhibit lower RHats ($\leq 1.05$) compared to GFSR ($\leq 1.25$), which may need more warm-up. The computation time of FAST is relatively robust to different $Q$ and $K$ values, so long as the chosen $K$ was at least the true value (see Supplement Section S4.4).
@@ 5 Data analysis: The DASH4D CGM Study
Dietary Approaches to Stop Hypertension for Diabetes (DASH4D, NCT04286555) is a nutritional trial designed to assess how blood pressure and glucose respond to combinations of DASH-style and low-sodium diets in persons with type 2 diabetes (T2D). The study team prepared four diets: DASH4D diet with lower sodium, DASH4D diet with higher sodium, Comparison diet with lower sodium, and Comparison diet with higher sodium. Weight was held constant by adjusting calorie level. Diet effects were observed through a single-site, 4-period crossover design (Pilla et al.,, 2025). Each period consisted of 5 weeks of feeding a diet followed by a $\geq 1$-week break (median 2 weeks). There were $N=105$ randomized T2D participants recruited from the Baltimore area, of which $N=65$ had meal timing data. Participants with meal timing data had a median age of $68$ years, were $66$% female, and had the following race distribution: $6.2$% Asian, $87.7$% Black, and $6.2$% White.
@@ 5.1 Mealtime CGM glucose
During feeding periods, participants ate meals at the study site 3 days each week. Clinic staff observed consenting participants to document meal timing. We used these timestamps to extract CGM data starting $1$ hour before and ending $4$ hours after each meal began. There were $\leq 20$ such $5-$hour periods per participant. The final dataset comprised 768 meals over 65 individuals. This data is not yet available as trial results are forthcoming.
Figure 3: Mealtime CGM for $4$ participants (one per row), for the Comparison (left column) and DASH4D (right column) diets. Sodium level within diet type is indicated by color. X-axis: time from meal initiation, y-axis: CGM in mg/dL. Person- and diet-specific average curves are darker, while meal-specific curves are lighter.
@@ 5.2 Bayesian FPCA of average CGM within diets
Figure 4: Bayesian FPCA results for the first three FPCs (columns) for each of the four diets (rows). X-axis: time relative to meal initiation. Black curves: posterior mean; red curves: three FPC posterior samples; shaded areas: point-wise $95$% credible intervals.
Figure 4 displays results for Bayesian FPCA inference for the first three PCs (columns) for the four diets (rows). For all panels the x-axis is the time from the start of the meal, the black curves are the posterior means, the red curves are three samples from the posterior of the PCs, and the shaded areas are point-wise, equal-tail $95$% credible intervals. The red curves correspond to the same sample iterations across PCs.
Figure 5: Eigenvalues (left panel) and percent-variability (right panel) estimates and $95$% credible intervals for each of the four diets (one color corresponds to one diet).
@@
Posterior eigenfunction estimates $\widehat{\phi}_{k}(t)$ are qualitatively similar across diets in Figure 4. The primary modes of variability $\widehat{\phi}_{1}(t)$ are nearly constant with a small curvature. Higher scores on these components correspond to higher CGM responses over the entire mealtime period. The sample-to-sample variability in the shape of $\phi_{1}(t)$ is low. Moreover, $\phi_{1}(t)$ explains $75$ to $85$% of the observed variability. This quantifies the importance of the mean postprandial glucose level (Garber,, 2012). The second FPCs, $\widehat{\phi}_{2}(t)$, generally exhibit an $S$ pattern. Therefore, individuals with a higher positive score on $\widehat{\phi}_{2}(t)$ have lower glucose in the pre-prandial and immediate postprandial period with a larger response one to four hours after eating. The third FPCs, $\widehat{\phi}_{3}(t)$, are negative during the hour before the meal, increase to a peak positive value about $50$ to $70$ minutes after food intake, and decrease back to negative values about $150-200$ minutes after the meal. Therefore, individuals with higher scores on $\widehat{\phi}_{3}(t)$ tend to have a to lower glucose before the meal and a rapid glucose increase after the meal. Sample-to-sample variability of both the $\phi_{2}(t)$ and $\phi_{3}(t)$ is higher than that of $\phi_{1}(t)$, likely due to the smaller signal.
The variability of glucose response in the DASH4D/lower sodium diet is much lower than for the other three diets. Figure 5 indicates that $\widehat{\lambda}_{1}(t)$ is $\approx 3$ times smaller for this diet compared to the others. This diet, the primary interest to the DASH4D study, appears to be associated with a more stable glycemic response between participants.
FAST produces reasonable results, qualitatively consistent with standard FPCA, across all four diets. Each model takes $20-45$ seconds (depending on the diet) to fit on a personal laptop using $5000$ iterations with the first $2500$ discarded as burn-in. We employ the routine from Section 2.5 to assess convergence of the scores, $\xi_{ik}$, and sampled points along the domain for each FPC, $\phi_{k}(t_{m})$. Final Gelman-Rubin statistics were $\leq 1.01$. Despite differences in sampling frequency per participant, results were similar when fitting FPCA to randomly sampled meals; see Supplement Section S5.2.

### v4

@@ 2.5 Alignment, Estimation, and Convergence
To monitor convergence of FAST we use the Gelman-Rubin RHat statistic for all model parameters (Gelman and Rubin,, 1992). For functional components, we have RHat values at all observation points. This is done after post-processing the PCs and scores.
@@ 3 Bayesian implementation in STAN
In this section, we provide the STAN implementation of FAST. STAN is a flexible probabilistic programming language which facilitates sampling from user-specified Bayesian models using Hamiltonian Monte Carlo (Betancourt,, 2018; Carpenter et al.,, 2017). We begin with the data section, which declares constants and imports the data, spline basis, and penalty matrix. We also calculate the trace of $\mathbf{P}_{\alpha}$ in the transformed data section for the FPC smoothing parameter priors. All notation is as used in Section 2.
@@
int M; // Number of observations
int Q; // Number of spline bases
int K; // Number of Eigenfunctions
matrix[N, M] Y; // Functional data
matrix[M, Q] B; // Orthogonal spline basis
matrix[Q, Q] P; // Penalty matrix for splines
transformed data {
real tr_P = trace(P_alpha); // Trace of penalty
In the parameters block, we model the matrix $\mathbf{X}$ (X). Through the parameter expansion described in Section 2.4, we indirectly model $\mathbf{\Psi}$ using the polar decomposition of $\mathbf{X}$. This block also contains the eigenvalues $\lambda_{k}$ (lambda) stored using a positive_ordered type, ordering the eigenvalues and corresponding eigenfunctions. The matrix Scores is the $N\times K$ score matrix $\mathbf{\Xi}$ introduced in Result 2.
@@
// Mean structure
vector[Q] w_mu; // Population mean parameters
real<lower=0> h_mu; // Population mean smoothing parameter
// Covariance structure
positive_ordered[K] lambda; // Eigenvalues
vector<lower=0>[K] H; // EF smoothing parameters
matrix[Q, K] X; // Unconstrained EF weights (X)
matrix[N, K] Scores; // EF scores (xi)
@@
// Polar decomposition
matrix[Q,K] Psi;
matrix[K,K] evec_XtX = eigenvectors_sym(X’*X);
vector[K] eval_XtX = eigenvalues_sym(X’*X);
Psi = X*evec_XtX*diag_matrix(1/sqrt(eval_XtX))*evec_XtX’;
The model proceeds with the specification of priors within the model section. The key components are the priors on $\mathbf{\Psi}$ (Psi). We first include the smoothing component directly through the penalties described in Section 2.3. Then, we incorporate the uniform component by specifying independent $N(0,1)$ priors for each entry of the latent matrix $\mathbf{X}$ (X). Combined with the polar decomposition in the transformed parameters block, this induces the uniform component of the prior on $\mathbf{\Psi}$. FAST thus models the FPCs through the small-dimensional $\mathbf{X},\boldsymbol{\Psi}\in\mathbb{R}^{Q\times K}$, both constrained by the size of the spline basis and number of FPCs. This dimensionality reduction likely contributes to the much improved computational efficiency. All other priors are as described in Section 2.3.
@@
H $\sim$ gamma(0.01, tr_P/2 + 0.01);
h_mu $\sim$ gamma(0.001, 0.001);
lambda $\sim$ inv_gamma(0.001, 0.001);
sigma2 $\sim$ inv_gamma(0.001, 0.001);
// Mean function smoothing prior
target += Q/2.0 * log(h_mu) - h_mu / 2.0 * w_mu * P * w_mu;
// FPC smoothing prior component
for(i in 1:K){
target += Q/2.0 * log(H[i]) - H[i] / 2.0 * Psi[,i]’ * P * Psi[,i];
// FPC uniform prior component
to_vector(X) $\sim$ normal(0,1);
// FPCA-based Score priors
for(i in 1:K){
to_vector(Scores[,k]) $\sim$ normal(0, sqrt(lambda[k]));
} $\cdots$
@@
model{ $\cdots$
{ // Data Likelihood
matrix[N,M] Theta = Scores * (B * Psi)’; // FPC linear combination
to_vector(Y) $\sim$ normal(to_vector(rep_matrix(mu, N) + Theta),
sqrt(sigma2));
@@ 4 Simulations
We consider the following two simulation scenarios for model (1), where $f_{k}(t)$ are taken to be the Legendre polynomials on $[0,1]$:
@@
For each scenario we generate $200$ datasets with $N=50$ functions observed at $M=50$ points distributed along $[0,1]$ according to quadrature. FAST is compared to the Generalized Function-on-Scalar Regression of Goldsmith et al., (2015) (labeled “GFSR”), modeling FPCs in the observed data space using parameter expansion based upon polar decomposition from Jauch et al., (2021) (labeled “POLAR”), and Variational Message Passing from Nolan et al., (2023) (labeled “VMP”). While all comparators fit a Bayesian FPCA, they do not all address uncertainty quantification for the same set of parameters. Neither POLAR nor VMP address construction of credible intervals for the functional components. Further, POLAR does not model the mean, instead subtracting point-wise means in pre-processing.
For the MCMC-based approaches, we performed 1500 sampling iterations and discarded the first 1000 as burn-in. This was chosen based upon the convergence properties of FAST and kept uniform between methods.
To obtain the orthonormal FPCs from the unconstrained samples of GFSR, we leverage the post-processing rotation suggested by Goldsmith et al., (2015). This involves calculating the SVD of the unconstrained FPC matrix sample and taking the right singular vectors. The scores are correspondingly transformed by the left singular vectors and diagonal matrix of singular values. This selection of a common rotation addresses the permutation- and rotation-based non-identifiability of the FPCs and scores. With a suitable alignment of sign, such as described in Section 2.5, the rotated posterior FPC/score samples can be used to form corresponding credible intervals.
Similar to FAST, both POLAR and GFSR require procedures to produce orthonormal (in the vector sense) FPC estimates. For POLAR, Jauch et al., (2021) recommends taking the right singular vectors of the mean smoothed data. As for GFSR, we take the point-wise mean of the non-orthogonal latent principal component matrix, then apply the SVD-based rotation Goldsmith et al., (2015) describes. VMP provides only FPC estimates.
POLAR, GFSR, and VMP all produce FPC estimates/samples which are orthonormal in the vector sense. So, we rescale the FPC estimates for these methods to have norm 1 in $L^{2}([0,1])$. As all methods are subject to the identifiability issues described in Section 2.5, we applied the post-processing steps described in that section to align posterior samples and estimates. FPC/score signs were then aligned with those of the true underlying FPCs for the purposes of evaluation.
@@
We also estimate pointwise posterior coverage probabilities of the true FPCs $\phi_{k}(t)$, $k=1,\ldots,K$. For each simulated data set $b\in\{1,\ldots,B\}$, we obtain the equal-tail 9595% credible intervals at the observed time points $t_{m}$, $m=1,\ldots,M$. The proportion of times these credible intervals cover the true function is calculated over the $t_{m}$ for each combination of $b$ and $k$. These proportions are then collected in a vector of length $B=200$ for each function $\phi_{k}(t)$. VMP and POLAR are excluded from this comparison, as they do not address uncertainty quantification for the FPCs.
Figure 2: Kernel smoother of 9595% credible interval coverage probabilities of the true FPCs for FAST, POLAR, and GFSR. Row indicates simulation scenario, S1 then S2. Each column corresponds to a particular eigenfunction. Distribution means: horizontal solid lines; nominal 9595% level: horizontal dotted lines.
Figure 2 displays the kernel smooth of the estimated point-wise coverage proportions for FAST and GFSR using the same structure as Figure 1. FAST performs well in all scenarios, consistently out-performing GFSR.
Figures 1 and  2 indicate that the advantage of FAST is particularly striking for $\phi_{3}(t)$ in S1. In this case, the mean can be written as a linear combination of $\phi_{1}(t)$ and $\phi_{3}(t)$: $\mu(t)=140\phi_{1}(t)+4\sqrt{5}\phi_{3}(t)$. This choice makes the problem more challenging in finite samples, when information may leak between the mean and covariance; see, for example, Zhou et al., (2025) for a more detailed discussion.
@@
We also compared estimation accuracy and posterior coverage of credible intervals for the scores $\xi_{k}$. Accuracy is calculated as mean square error aggregated by FPC, $k$, within each simulation, $b$. Coverage was calculated as the proportion of times the equal-tail 9595% credible intervals cover the corresponding true scores; see results in Supplement Section S4.3. FAST outperforms the other methods and has near nominal coverage.
@@
$M$ FAST GFSR POLAR VMP NN FAST GFSR POLAR VMP
50 $3.5$ $24.3$ $9.2$ $0.2$ 50 $3.5$ $24.3$ $9.2$ $0.2$
100 $6.8$ $39.5$ $28.5$ $1.0$ 100 $4.3$ $43.2$ $15.4$ $0.6$
250 $18.3$ $82.8$ $1.2$ 250 $5.3$ $47.2$ $29.0$ $0.6$
500 $37.6$ $151.0$ $5.2$ 500 $14.7$ $27.8$ $52.1$ $0.7$
@@
FAST and GFSR scale approximately linearly with $M$, though GFSR has a substantially higher slope. FAST and POLAR scale linearly in $N$, while GFSR seems to be more performant with higher $N$, possibly due to the increased information better specifying the posterior geometry. We could not fit POLAR for $M\geq 250$ because it uses too much memory. VMP is the fastest, which is not unexpected given the approximation inherent to the technique. All MCMC-based methods were run for 1500 iterations with the first 1000 discarded as burn-in; however, the convergence characteristics were not uniform. FAST and POLAR exhibit lower RHats ($\leq 1.05$) compared to GFSR ($\leq 1.25$), which may require a longer warm-up. The computation time of FAST is relatively robust to different $Q$ and $K$ values (see Supplement Section S5.4).
@@ 5 Data analysis: The DASH4D CGM Study
Dietary Approaches to Stop Hypertension for Diabetes (DASH4D, NCT04286555) is a nutritional trial designed to assess how blood pressure and glucose respond to combinations of DASH-style and low-sodium diets in persons with type 2 diabetes (T2D). The study team prepared four diets: DASH4D diet with lower sodium, DASH4D diet with higher sodium, Comparison diet with lower sodium, and Comparison diet with higher sodium. Weight was held constant by adjusting calorie level. Diet effects were observed through a single-site, 4-period crossover design (Pilla et al.,, 2025). Each period consisted of 5 weeks of feeding a diet followed by a $\geq 1$-week break (median 2 weeks). Approval for DASH4D was granted by the Johns Hopkins University School of Medicine institutional review board, and all study participants provided informed written consent. There were $N=105$ randomized T2D participants recruited from the Baltimore area, of which $N=65$ had meal timing data. Participants with meal timing data had a median age of $68$ years, were 6666% female, and had the following race distribution: 6.26.2% Asian, 87.787.7% Black, and 6.26.2% White.
@@ 5.1 Mealtime CGM glucose
During feeding periods, participants ate meals at the study site 3 days each week. Clinic staff observed consenting participants to document meal timing. We used these timestamps to extract CGM data starting $1$ hour before and ending $4$ hours after each meal began. There were $\leq 20$ such 55-hour periods per participant. The final dataset was comprised of 768 meals over 65 individuals. This data is not yet available as trial results are forthcoming.
Figure 3: Mealtime CGM for 44 participants (one per row), for the Comparison (left column) and DASH4D (right column) diets. Sodium level within diet type is indicated by color. X-axis: time from meal initiation, y-axis: CGM in mg/dL. Person- and diet-specific average curves are darker, while meal-specific curves are lighter.
@@ 5.2 Bayesian FPCA of average CGM within diets
Figure 4: Bayesian FPCA results for the first three FPCs (columns) for each of the four diets (rows). X-axis: time relative to meal initiation. Black curves: posterior mean; red curves: three FPC posterior samples; shaded areas: point-wise 9595% credible intervals.
Figure 4 displays results for Bayesian FPCA inference for the first three PCs (columns) for the four diets (rows). For all panels the x-axis is the time from the start of the meal, the black curves are the posterior means, the red curves are three samples from the posterior of the PCs, and the shaded areas are point-wise, equal-tail 9595% credible intervals. The red curves correspond to the same sample iterations across PCs.
Figure 5: Eigenvalues (left panel) and percent-variability (right panel) estimates and 9595% credible intervals for each of the four diets (one color corresponds to one diet).
@@
Posterior eigenfunction estimates $\widehat{\phi}_{k}(t)$ are qualitatively similar across diets in Figure 4. The primary modes of variability $\widehat{\phi}_{1}(t)$ are nearly constant with a small curvature. Higher scores on these components correspond to higher CGM responses over the entire mealtime period. The sample-to-sample variability in the shape of $\phi_{1}(t)$ is low. Moreover, $\phi_{1}(t)$ explains $70$ to 8585% of the observed variability. This quantifies the importance of the mean postprandial glucose level (Garber,, 2012). The second FPCs, $\widehat{\phi}_{2}(t)$, generally exhibit an $S$ pattern. Therefore, individuals with a higher positive score on $\widehat{\phi}_{2}(t)$ have lower glucose in the pre-prandial and immediate postprandial period with a larger response one to four hours after eating. The third FPCs, $\widehat{\phi}_{3}(t)$, are negative during the hour before the meal, increase to a peak positive value about $50$ to $70$ minutes after food intake, and decrease back to negative values about 150200150-200 minutes after the meal. Therefore, individuals with higher scores on $\widehat{\phi}_{3}(t)$ tend to have a to lower glucose before the meal and a rapid glucose increase after the meal. Sample-to-sample variability of both the $\phi_{2}(t)$ and $\phi_{3}(t)$ is higher than that of $\phi_{1}(t)$, likely due to the smaller signal.
The variability of glucose response in the DASH4D/lower sodium diet is much lower than for the other three diets. Figure 5 indicates that $\widehat{\lambda}_{1}(t)$ is $\approx 3$ times smaller for this diet compared to the others. This diet, a primary interest of the DASH4D study, appears to be associated with a more stable glycemic response between participants.
FAST produces reasonable results, qualitatively consistent with standard FPCA, across all four diets. Each model takes 104010-40 seconds (depending on the diet) to fit on a personal laptop using $2000$ iterations with the first $1000$ discarded as burn-in. We employ the routine from Section 2.5 to assess convergence of the scores, $\xi_{ik}$, and sampled points along the domain for each FPC, $\phi_{k}(t_{m})$. Final Gelman-Rubin statistics were <1.05<1.05. Despite differences in number of meals per participant, results were similar when fitting FPCA to randomly sampled meals; see Supplement Section S6.2.